# Project Page: https://yikai-wang.github.io/asuka/ MISATO Dataset: https://github.com/Yikai-Wang/asuka-misato

### v2

@@ 3.2 Align MAE Prior with SD for Context-Stability
Replace text-condition with MAE prior. SD is not trained on MAE priors. Fortunately, the text-condition part of SD is not utilized in inpainting, as we do not assume a text condition for inpainting task. Hence, we propose to replace the text-condition part of SD with our proposed MAE prior condition. However, as we do not fine-tune the SD, it cannot aligns well with the MAE prior. To achieve a proper guidance for SD, we introduce the alignment module to align MAE with SD in both dimension and distribution perspective, as shown in 3(b).
@@
Dimension alignment. Particularly, the MAE prior $F_{\mathrm{MAE}}$ is of size $N_{m}\times M_{m}$, where $N_{m}$ is the sequence length and $M_{m}$ is the feature dimension. To align it with the SD condition of size $N_{s}\times M_{s}$, we adopt a linear layer to map the feature dimension from $M_{m}$ to $M_{s}$ and set $N_{s}=N_{m}$ to preserve the local MAE prior.
@@
Distribution alignment. After aligning the dimension, we use self-attention blocks to learn to better guiding SD, leading to the condition $C_{\mathrm{MAE}}$ for SD. We train our alignment module using the standard diffusion objective with the same masking strategy used to train the MAE, keeping other modules frozen.
@@
Misalignment. When training the alignment module with the set (input image, MAE prior, SD result), misalignment issues may arise. For example, if an object is completely masked, the MAE will predict the masked area with background, whereas the SD is trained to recreate the object. This difference can lead the alignment module to mistakenly disregard the MAE prior. To address this, we improve the SD’s adherence to the MAE prior by substituting the MAE predicted prior with the MAE reconstructed prior at a probability of $p$. The MAE reconstructed prior involves using MAE to recreate the image without masking any area, ensuring MAE has access to all information needed for reconstruction. This approach helps train the alignment module to better guide the SD.
Figure 5: The illustration of the information loss of KL-VAE. (a) The hue of the image is shifted. (b) KL-VAE still suffers from non-ignorable shifts in low-frequency fields.
(a) Qualitative SD1.5 inpainting results decoded by (b) vanilla decoder of SD [39], (c) larger decoder [62], (d) our decoder. Our decoder largely alleviate the mask-unmask color inconsistency.
(b) Decoder trained by local harmonization task, enhancing mask-unmask consistency by reconstructing original image guided by the unmasked region from augments in color and latent spaces.
Figure 6: ASUKA trains a specialized-decoder to solve the color-inconsistency issue from the information loss of KL-VAE and domain gap between generated and real latent.
3.3 Mask-Unmask Align during Decoding for Visual-Consistency
@@
Information loss of KL-VAE. SD performs all its generative processes in the latent space and subsequently decodes these latent codes back to image space using KL-VAE. Despite the decoder being trained to reconstruct the image, it encounters challenges associated with information loss. Particularly in tasks like inpainting, we have ground-truth values for the unmasked region. Though Rombach et al. [39] claimed that the diffusion model should prioritize the informative semantic compression, while the VAE and GAN are used to tackle perceptual compression with high-frequency details, we argue that low-frequency semantic loss in KL-VAE could not be neglected, as verified in Fig. 5 (b). The KL-VAE will not only noticeably degrades high-frequency details but also shifts in semantic hue. This shift can be verified by repeated reconstruction with KL-VAE, as shown in Fig. 5 (a) where larger shift is observed during repeated reconstruction. As human is sensitive to low-frequency information changes in the image, even subtle color shifts can induce significant visual-inconsistencies. This visual-inconsistency issue is more severe under the irregular or large mask structures.
@@
Unmask-region conditioned decoder. The aforementioned information loss is prohibitive to be addressed by the SD U-Net alone, especially given the already degraded input latent codes. Zhu et al. [62] adopts a larger decoder, with additional inputs of masked images. However, it still fails to handle the incompatible color and texture between the original images and compressed ones in challenging scenes as verified in Fig. 5(a) (c). The low-frequency gap between degraded and original images makes it challenging to explicitly address this issue. To solve the mask-unmask visual-inconsistency issue, we reformulate the decoding as a local harmonization task to preserve the visual consistency. As shown in Fig. 3, compared with SD decoder, our decoder involves additional inputs of masked images in the pixel-wise color space and the 0-1 mask.
@@
Mask-unmask visual-consistent decoder. To train the decoder to ensure visual-consistency between generated latent and unmasked pixels, we re-formulate the decoding as a local harmonization task. Specifically, we propose the color and latent augmentation as shown in Fig. 5(b) to estimate and enlarge the visual-inconsistency during decoding. We follow the standard KL-VAE training pipeline of SD, but replacing the inputs with augmented ones. Particularly, we use the original image as the reconstruction target and use color and latent augmentation to corrupt input image, simulating the information loss of KL-VAE and domain gap between generated and real latent, respectively. This forces the decoder to reconstruct the un-corrupted image based on the ground-truth unmasked region.
@@
i) Color augmentation: The first information loss comes from the imperfect VAE reconstruction. VAE is trained to encode and decode the input image for reconstruction, but suffers from color shift as shown in Fig. 5(a) (b). Empirically, further conditioned on unmasked image alleviate but not solve the color inconsistency issue, as shown in Fig. 5(a) (c). Hence, we need to explicitly train the decoder to ensure color consistency. To this end, we augment all training images in brightness, contrast, saturation, and hue, and requires the decoder to reconstruct original image conditioned on the unaugmented unmasked image. This encourages the decoder to fully utilize the information from unmasked regions.
@@
ii) Latent augmentation: The other information loss comes from the distribution gap between SD generated latent and real latent. This gap also causes color inconsistency even if we alleviate the VAE reconstruction loss, as shown in Fig. 6(a). To simulate this loss, we incorporate the artifacts generated from the SD to train the decoder. However, denoising to real images iteratively is notably time-consuming, even with DDIM [42]. To balance the efficiency and efficacy, we design a one-step estimation of SD. As our target is to capture the generation gap, we use the clean latent $\bm{z}_{0}$ and all-zero mask $\mathbf{M}$ as conditions. This tells the SD all the needed information to generate the clean latent, ensuring the generated latent preserves content and only shift from the distribution gap. We follow the standard pipeline to estimate $\bm{z}_{0}$ with modified conditions as:
$\hat{\bm{z}}_{0}=\frac{1}{\sqrt{\overline{\alpha}_{t}}}(\bm{z}_{t}-\sqrt{1-\overline{\alpha}_{t}}\varepsilon_{\theta}\left([\bm{z}_{t};\bm{z}_{0};\mathbf{M}],t\right)),$ (1)
@@
where the timestep $t$ is randomly sampled from $[500,1000)$; $\overline{\alpha}_{t}$ indicates the prescribed variance schedule in SD; $\varepsilon_{\theta}(\cdot)$ is the frozen SD take as inputs noised $\bm{z}_{t}$, unmasked $\bm{z}_{0}$, and all-zero masking $\mathbf{M}$. The large step denoising is chosen to increase the distribution gap, as empirically SD could produce reliable results in small $t$ given the unmasked latent condition $\bm{z}_{0}$. Then we decode $\hat{\bm{z}}_{0}$ to image as the latent augmented inputs. This makes latent augmentation an off-line strategy. We apply latent augmentation to 50% training images. We fine-tune the decoder based on [62], showcasing superior consistency as compared in Fig. 5(a).
(a) w/ v.s. w/o latent aug.
(b) Different image domains in MISATO.
Figure 7: (a) Ablation of latent augmentation; (b) MISATO evaluation examples.
4 Evaluation Dataset
@@
In addition to the standard benchmark Places 2 [61], to validate the inpainting performance across different domains and mask styles, we construct a evaluation dataset, dubbed as MISATO, from Matterport3D [10], Flickr-Landscape [28], MegaDept[26], and COCO 2014 [29]. The first three datasets contains images from indoor, outdoor landscape and building, respectively. COCO has object segmentation. We select 500 representative examples of size $512^{2}$ from each dataset. The MISATO dataset is available at https://github.com/Yikai-Wang/asuka-misato.

### v3

@@ 3.1.2 Align MAE Prior with Generator
Generative inpainting models are not trained on MAE priors. As we do not assume a text condition for inpainting task, we propose to replace the text-condition of generative models with our proposed MAE prior condition. However, as we do not fine-tune the generative models, they cannot directly align well with the MAE prior. Hence, we introduce the alignment module to align MAE with generative models in both dimension and distribution perspective, as shown in Fig. 2(b).
Dimension alignment
@@
Particularly, the MAE prior $F_{\mathrm{MAE}}$ is of size $N_{m}\times M_{m}$, where $N_{m}$ is the sequence length and $M_{m}$ is the feature dimension. To align it with the diffusion or flow condition of size $N_{s}\times M_{s}$, we adopt a linear layer to map the feature dimension from $M_{m}$ to $M_{s}$ and set $N_{s}=N_{m}$ to preserve the local MAE prior.
Distribution alignment
@@
After aligning the dimension, we use self-attention blocks to learn to better guiding generative models, leading to the condition $C_{\mathrm{MAE}}$. We train our alignment module using the standard generation objective with the same masking strategy used to train the MAE, keeping other modules frozen.
Handle misalignment
@@
When training the alignment module with the set (input image, MAE prior, inpaint result), misalignment may arise. For example, if an object is completely masked, the MAE will predict the masked area with background, whereas the generative models are trained to recreate the object. This difference can lead the alignment module to mistakenly disregard the MAE prior. To address this, we improve the generative models’ adherence to the MAE prior by substituting the MAE predicted prior with the MAE reconstructed prior at a probability of $p$. The MAE reconstructed prior involves using MAE to recreate the image without masking any area, ensuring MAE has access to all information needed for reconstruction. This approach helps train the alignment module to better guidance.
3.2 Enhancing Color-Consistency in Decoding
3.2.1 Color-Inconsistency
Color-inconsistency is a general problem
The color-inconsistency between masked and unmasked regions is a general problem in generative inpainting models. This inconsistency comes when the generative masked region suffers from a color shift compared with the unmasked region. As in Fig. 4, the color shift happens in all kinds of scenarios, including indoor and outdoor scenes, random or continuous masks, and may cause darker or lighter color shift. This shift comes from the imperfect VAE and latent generator.
Information loss of VAE
@@ 3.2.1 Color-Inconsistency
Popular latent diffusion and rectified flow models perform all the generative processes in the latent space and subsequently decodes these latent codes back to image space using VAE. Despite the decoder being trained to reconstruct the image, it encounters challenges associated with information loss. Particularly in tasks like inpainting, we have ground-truth values for the unmasked region. Though Rombach et. al. [67] claimed that the diffusion model should prioritize the informative semantic compression, while the VAE is used to tackle perceptual compression with high-frequency details, we argue that low-frequency semantic loss in VAE could not be neglected, as verified in Fig. 6 (b). The VAE will not only noticeably degrades high-frequency details but also shifts in colors. This shift can be verified by repeated reconstruction with VAE, as shown in Fig. 6 (a) where larger shift is observed during repeated reconstruction. As human is sensitive to low-frequency information changes in the image, even subtle color shifts can induce significant inconsistencies. This issue is more severe in irregular or large mask cases.
Figure 4: The color shift exists in all kinds of scenarios in inpainted images, including indoor and outdoor scenes, random or continuous masks, and may cause darker or lighter color shift.
Figure 5: Inpainting w/ v.s. w/o latent augmentation. The latent augmentation handles the gap between generated and real latent.
Gap between real and generated latents
Apart from the information loss of VAE in reconstruction, there is another gap between the generated and real latents. This gap also causes color inconsistency even if we alleviate the VAE reconstruction loss, as in Fig. 5. We need to solve both the loss of VAE and the latent generator for better color-consistency.
Figure 6: (a) The color of the reconstruted image is shifted, where larger shift is observed during repeated reconstruction. (b) VAE suffers from non-ignorable shifts in low-frequency fields.
Figure 7: Decoder trained by local harmonization task, enhancing mask-unmask consistency by reconstructing original image guided by the unmasked region from augments in color and latent spaces.
Figure 8: SD1.5 inpainting results decoded by (b) vanilla decoder of SD [67], (c) conditional decoder [103], (d) our decoder. Our decoder largely alleviate the mask-unmask color inconsistency.
3.2.2 Mask-Unmask Align during Decoding
We propose to solve the color-inconsistency and ensure the mask-unmask alignment during VAE decoding.
Unmask-region conditioned decoder
@@ 3.2.2 Mask-Unmask Align during Decoding
The basic solution is to incorporate the ground-truth unmasked region in the decoding, then we could have access to the unbiased color information. Zhu et. al. [103] adopts decoder with additional inputs of masked images. However, it still fails to handle the incompatible color and texture between the original images and compressed ones in challenging scenes as verified in Fig. 8 (c). The gap between degraded and original images makes it challenging to explicitly address this issue.
Mask-unmask color-consistent decoder
@@
To train the decoder to ensure color-consistency between generated latent and unmasked pixels, we re-formulate the decoding as a local harmonization task. Our decoder involves additional inputs of masked images in the pixel-wise color space and the 0-1 mask. To properly train the decoder, we propose the color and latent augmentation as shown in Fig. 7 to estimate and enlarge the color-inconsistency. We follow the standard VAE training pipeline, but replacing the inputs with augmented ones. Particularly, we use the original image as the reconstruction target and use color and latent augmentation to corrupt input image, simulating the information loss of VAE and domain gap between generated and real latent, respectively. This forces the decoder to reconstruct the clean image based on the ground-truth unmasked region.
Color augmentation
@@
We use color augmentation to capture the VAE loss as in Fig. 8 (b). Empirically, further conditioned on unmasked image alleviate but not solve the color inconsistency issue, as shown in Fig. 8 (c). Hence, we need to explicitly train the decoder to ensure color consistency. To this end, we augment all training images in brightness, contrast, saturation, and hue, and requires the decoder to reconstruct original image conditioned on the unaugmented unmasked image. This encourages the decoder to faithfully follow the unmasked regions.
Latent augmentation
@@
To simulate the gap between generated and real latent, we incorporate the artifacts generated from the generative models to train the decoder. However, denoising to real images iteratively is notably time-consuming, even with DDIM [73]. To balance the efficiency and efficacy, we design a one-step estimation. As our target is to capture the generation gap, we use the clean latent $\bm{z}_{0}$ and all-zero mask $\mathbf{M}$ as conditions. This tells the generator all the needed information to generate the clean latent, ensuring the generated latent preserves content and only shift from the generation gap. We follow the standard pipeline to estimate $\bm{z}_{0}$ with modified conditions as:
$\hat{\bm{z}}_{0}=\frac{1}{a}(\bm{z}_{t}-b\varepsilon_{\theta}\left([\bm{z}_{t};\bm{z}_{0};\mathbf{M}],t\right)),$ (1)
@@
where the timestep $t$ is randomly sampled from $[500,1000)$; $a$ indicates the prescribed variance schedule, $a^{2}+b^{2}=1$ in diffusion models while $a+b=1$ in rectified flow models; $\varepsilon_{\theta}(\cdot)$ is the frozen generator take as inputs noised $\bm{z}_{t}$, unmasked $\bm{z}_{0}$, and all-zero masking $\mathbf{M}$. The large step denoising is chosen to increase the distribution gap, as empirically the generator could produce reliable results in small $t$ given the unmasked latent condition $\bm{z}_{0}$. Then we decode $\hat{\bm{z}}_{0}$ to image as the latent augmented inputs. This makes latent augmentation an off-line strategy. We apply latent augmentation to 50% training images. The fine-tuned decoder showcases superior consistency as compared in Fig. 8.
Figure 9: Inpainting results for 5122 images. GANs generate blurred results; SD variants hallucinate unreasonable objects and suffer from color shift. ASUKA achieves unwanted-object-mitigated and color-consistent inpainting. More results are in the appendix.
4 Experiments
Evaluation datasets
We follow previous works to evaluate on the standard benchmark Places 2 [101] validation set of 36,500 images. In addition, to validate across different domains and mask styles, we construct a evaluation dataset, dubbed as MISATO, from Matterport3D [13], Flickr-Landscape [47], MegaDept[45], and COCO 2014 [48] to handle indoor, outdoor, building and background inpainting, respectively. We select 500 representative examples of size $512^{2}$ and $1024^{2}$ from each dataset, forming a total of 2,000 testing examples. See details in the appendix.
General evaluation metrics
We use the Learned Perceptual Image Patch Similarity (LPIPS) [97] to calculate the patch-level image distances, Fréchet Inception Distance (FID) [32] to compare the distribution distance between generated images and real images, and Paired/Unpaired Inception Discriminative Score (P-IDS/U-IDS) [99] to measure the human-inspired linear separability.
Evaluate object hallucination and color-consistency
We introduce two new metrics to assess the object hallucination and color-consistency of inpainted images. (1) CLIP@mask (C@m): We use CLIP to get visual features from both the ground-truth and the inpainted masked region, then calculate their cosine similarity. Following the standard CLIP score, we multiply the result by 100 and clip negative values, yielding a range from 0 to 100. (2) Gradient@edge (G@e): We calculate the average pixel gradient difference along the edges of the masked region with respect to the ground-truth image to assess color smoothness. A smaller gradient difference means more similar color transitions to the ground-truth image and, therefore, less color shift.
Table 1: Quantitative comparison on MISATO and Places 2. Top-3 results are colored.
Dataset MISATO Places 2
@@ 4 Experiments
Co-Mod [99] 0.179 17.421 0.243 0.109 0.924 52.106 0.267 5.794 0.274 0.096 0.951 166.914
MAT [44] 0.176 17.261 0.255 0.122 0.925 48.722 0.202 3.765 0.348 0.195 0.955 163.442
LaMa [75] 0.155 15.436 0.260 0.135 0.928 46.270 0.202 6.693 0.247 0.050 0.953 153.653
MAE-FAR [10] 0.142 13.283 0.282 0.153 0.940 43.613 0.174 3.559 0.307 0.105 0.958 149.843
SD-Repaint [54] 0.227 27.861 0.016 0.007 0.915 80.410 0.251 12.466 0.217 0.045 0.947 176.421
SD [67] 0.168 12.812 0.345 0.211 0.951 63.844 0.193 1.514 0.375 0.207 0.959 160.705
SD-text 0.164 12.603 0.337 0.207 0.952 63.776 0.191 1.506 0.373 0.202 0.959 160.418
SD-token [81] 0.160 12.517 0.331 0.204 0.955 61.700 0.189 1.477 0.390 0.234 0.960 158.924
SD-IP [88] 0.157 12.204 0.398 0.242 0.956 62.704 0.186 1.539 0.389 0.173 0.953 148.571
SD-T2I [58] 0.166 13.806 0.365 0.222 0.949 63.866 0.195 1.720 0.384 0.160 0.951 148.549
SD-CAEv2 [98] 0.157 29.179 0.193 0.045 0.901 69.890 0.192 6.887 0.287 0.065 0.921 151.863
SD-LaMa [75] 0.157 12.159 0.390 0.256 0.956 62.726 0.188 1.522 0.389 0.168 0.953 148.461
ASUKA-SD 0.150 11.495 0.423 0.312 0.958 47.753 0.183 1.230 0.413 0.287 0.961 147.733
Competitors
We primarily use the SD v1.5 inpainting model [67] to analyze and compare ASUKA with competitors, while validating ASUKA’s generalization ability with FLUX. We consider three SD v1.5 inpainting variants: SD: uses a null-prompt for unconditional generation; SD-text: uses "background" as a prompt since no captions are used in inpainting; SD-token [81]: uses learnable tokens trained with ASUKA’s pipeline. To test other ways of incorporating the MAE condition, we implement the following: SD-IP, uses IP-Adapter [88]; SD-T2I, uses T2I-Adapter [58]; SD-CAEv2, uses a CLIP-style alignment module CAEv2 [98]; We also test SD-LaMa, which inputs LaMa [75] inpainting results instead of MAE. We also compare with leading inpainting algorithms Co-Mod [99], MAT [44], LaMa [75], MAE-FAR [10], and SD-Repaint [54]. We provide implementation details in the appendix.
4.1 Comparison on Benchmarks
Quantitative comparison
Results on SD are reported in Tab. 1. (1): Although ASUKA-SD is based on a fixed SD model, it consistently outperforms SD across all evaluation metrics, achieving state-of-the-art results in FID, U-IDS, and P-IDS. Notably, U-IDS and P-IDS are closely aligned with human preferences [99] and have a potential maximum score of 0.5, highlighting ASUKA’s strong performance. (2): Compared to other adapters that align the MAE prior with SD, ASUKA-SD shows consistently superior performance across all metrics. This demonstrates the effectiveness of our straightforward alignment module. (3): While the LaMa condition improves inpainting quality, as shown by FID and IDS scores, it is less effective than the MAE condition. When using the MAE condition as a prior, improvements can be attributed to better mitigation of object hallucination. (4): ASUKA-SD consistently performs better than all competitors on CLIP@mask, showcasing the strength of its improved mitigation of object hallucination. (5): Pixel-based GAN inpainting models perform better in the Gradient@edge metric, suggesting that color shifts may originate from the compressed latent space. ASUKA-SD, however, still shows significant improvements over all SD variants, highlighting its enhanced color consistency. (6): The second-to-best LPIPS scores are partially due to using a frozen SD, where ASUKA achieves consistent improvements but remains constrained by the frozen U-Net. These results confirm that ASUKA-SD improves color consistency and mitigation of object hallucination in inpainting, even when using frozen latent inpainting models. This advantage is evident both in the in-distribution dataset Places2 and the out-of-distribution dataset MISATO.
Qualitative comparison
examples are shown in Fig. 9. (1) The state-of-the-art inpainting algorithms usually suffer from unnatural generation, for example the unnatural boundaries in the third and fourth rows, and failed inpainting of tower in the third-to-last row. LaMa and MAE-FAR sometimes lead to blurred inpainting results, especially in the scenario of large continuous masks. (2) The SD variants usually suffer from the unwanted object insertion issue and hallucinate unreasonable objects, in almost all the illustrated images. (3) In contrast, ASUKA enjoys unwanted-object-mitigated and color-consistent inpainting.
4.2 Further Analysis of ASUKA
In this part, we conduct more experiments to analysis ASUKA. More analysis can be found in the appendix.
Extension to FLUX
To demonstrate ASUKA’s versatility, we trained it on FLUX (see Tab. 2). ASUKA-FLUX consistently outperforms the original FLUX. Results on Places 2 and qualitative comparisons are in the appendix.
Table 2: FLUX and ASUKA-FLUX on MISATO@512. Results on 1K and qualtitative results are in the appendix.
[…]
FLUX 0.254 12.839 0.351 0.223 0.951 65.928
ASUKA-FLUX 0.206 11.372 0.428 0.327 0.962 48.635
Table 3: Comparison of different decoders for SD.
[…]
VAE 0.156 11.949 0.387 0.253 0.953 63.142
+ cond. 0.151 11.634 0.410 0.272 0.955 48.588
+ color 0.152 11.603 0.407 0.273 0.954 49.538
Ours 0.150 11.495 0.423 0.312 0.958 47.753
Ablation of decoder
For the decoder, we compare ASUKA-SD with (1) VAE: the decoder used in SD; (2) + cond.: the decoder conditioned on unmasked image [103]; (3) + color: only trained with color augmentation ; Results are in Tab. 3, showing the superiority of our decoder.
Table 4: Ablation of different alignment modules.
[…]
linear 0.155 11.934 0.400 0.263 0.953 48.983
attn 0.152 11.613 0.403 0.268 0.954 48.785
cross x4 0.152 11.762 0.405 0.256 0.953 48.279
Ours 0.150 11.495 0.423 0.312 0.958 47.753
Ablation of alignment module
We validate the efficacy of our alignment module step by step: (1) linear: Use linear layer to align feature dimension only; (2) attn: Based on linear, further use a single self-attention block to align the distribution; (3) cross x4: we instead use learnable query and 4 cross-attention layers to learn the MAE prior. ASUKA-SD adopts 4 self-attention blocks. Results are shown in Tab. 4. The self-attention block shows improved results compared with only align dimension and cross-attention block. Using 4 self-attention blocks improves the capacity.
5 Conclusion
In this paper, we proposed Aligned Stable inpainting with Unknown Areas prior (ASUKA) to achieve unwanted-object-mitigated and color-consistent inpainting via frozen latent inpainting models. To avoid unwanted object insertion, we adopt a reconstruction-based masked auto-encoder (MAE) as the context-stable prior for masked region purely from unmasked region. Then we align the context-stable prior to frozen generative models with the proposed alignment module. To achieve color-consistency, we resolve the mask-unmask color inconsistency in the latent decoding process. We train an unmask-region conditioned VAE decoder to perform local harmonization during the decoding process. To validate the efficacy of inpainting algorithms in different image domains and mask types, we introduce an evaluation dataset, named as MISATO, from existing datasets. We propose two new metrics to explicitly evaluate the object hallucination and color-consistency of inpainted images. ASUKA enjoys unwanted-object-mitigated and color-consistent inpainting results and superior than leading inpainting models.
Acknowledgments
The authors would like to thank Huawei Ascend Cloud Ecological Development Project for the support of Ascend 910 processors.
Supplementary Material
6 Brief Introduction of Backbone Models
We evaluate our proposed solution on two inpainting models: the Stable Diffusion v1.5 inpainting model (SD) [67] and the Control-Net fine-tuned FLUX inpainting model (FLUX) [2]. Both models are representative latent inpainting models that use a VAE [38] to compress images into a smaller latent space. In SD, a diffusion process [72] maps the latent space to random Gaussian noise, and a U-Net [68] learns the reverse denoising path. Text condition is introduced through cross-attention layers [77]. The inpainting version of SD extends the U-Net input by concatenating the masked image and mask with the noise along the channel dimension. Conversely, FLUX uses rectified flow [52, 1, 49] to map the latent space to noise and a vision transformer [63] for generation. Text condition is applied by concatenating text with image patches as transformer input, while a pooled text condition is injected into the normalization layers. Since the original FLUX [40] does not support inpainting, we use a Control-Net [96] fine-tuned version [2] that modifies FLUX’s transformer output by conditioning on the masked image and mask. We demonstrate that our ASUKA effectively improves unwanted object mitigation and color consistency of these models.
7 Details about MISATO